OpenStudy (anonymous):

Factor the expression: x^2 - 7x + 10

OpenStudy (anonymous):

(x-2)(X-5)

OpenStudy (math&ing001):

We can do it like we're gonna solve the quadratic equation x² - 7x + 10 = 0. @alexiselizabeth we value the guiding part more than giving out answers.

OpenStudy (anonymous):

I put that and I got it wrong.

OpenStudy (math&ing001):

@OM14forever do you know how to solve quadratic equations ?

OpenStudy (anonymous):

(X+2) (x+5)

OpenStudy (anonymous):

I keep getting 2 and 5

OpenStudy (math&ing001):

Show me how you proceeded, maybe I can detect where you made the mistake.

OpenStudy (anonymous):

X^2-7x+10 (X-2)(x+5)

OpenStudy (anonymous):

I got 2 and 5 because they are factor of 10 and when you add them they make seven.

OpenStudy (math&ing001):

If you want to make sure you got your factoring correct try distributing it and see if it matches with your first expression. Here (x-2)(x+5)= x² +3x -10 which is not correct. Here's a way of doing it: x² - 7x + 10 = x² - 2x - 5x + 10 = x(x-2) - 5(x-2) = (x-5)(x-2)